Problem Seven The table lists drive-thru order accuracy at popular fast food chains (data from a 2016 QSR Drive-Thru Study). Assume that orders are randomly selected from those included in the table.

McDonalds 334 Burger King 288 Wendys 243 Taco Bell 145 Order Accurate Order Not Accurate 28 30 37 13

If one order is selected, what is the probability of the order:

A. Being Accurate?

B. NOT being from Wendy’s?  

C. Being from McDonald’s OR NOT Accurate?

D. Are the events of selecting an order from McDonald’s and selecting an accurate order mutually exclusive events? Explain your reasoning.
Answer #1
McD Burger King Wendy's Taco bell Total
Accurate 334 288 243 145 1010
not accurate 28 30 37 13 108
Total 362 318 280 158 1118

A)

P( Being Accurate ) =1010/1118 =0.9034

B)

P( NOT being from Wendy’s )=1-P(from Wendy’s ) =1-280/1118 = 0.7496

C)

P( Being from McDonald’s OR NOT Accurate ) =P(McD)+P(not Accurate)-P(MCD and Not Accurate)

=(362/1118)+(108/1118)-(28/1118)

=0.3953

D)

No since both events contains 334 element in common